Mortimer lives for two period and has utility function U = C1*C2. He earns no income in period two and his income in period 1 is $80,000. The interest rate at which he can borrow and lend is 10%. Calculate his optimal consumption in each period.
Answer #1

Mortimer's budget constraint= C1+(1/(1+.1)*C2 =80000

MRS=C2/C1=1/( 1/1.1)= P1/P2

C2=1.1*C1

Put this equation in budget constraint

C1+1.1C1/1.1=80000

2C1= 80000

C1*= $40000

C2=1.1*40000= $44000

Optimal consumption in period 1=$40000

Optimal consumption in period 2= $44000